OpenStudy (anonymous):

What is the value of xy over z if x = –1, y= –10 and z = 2? 5 –5 one-fifth –one-fifth

OpenStudy (anonymous):

\(x = -1\), \(y = -10\), \(z = 2\) and : Find \(\large \frac{xy}{z}\), we are to just place the values of \(x,y,z\) in their places: \[\frac{x \times y}{z} = \frac{-1 \times -10}{2} = ??\]

OpenStudy (anonymous):

Firstly tell : \(-10 \times -1 = ??\)

OpenStudy (anonymous):

I saw you closing your question.. :P